What is the wavelength of an electromagnetic radiation that ejects 2.00-ev electrons from calcium metal,given that the binding energy is 2.71ev,and what type of electromagnetic radiation is this?
1
Expert's answer
2021-08-20T18:04:54-0400

"h\frac c{\lambda}=E_2-E_1,"

"\lambda=\frac{hc}{E_2-E_1}=1.75\cdot 10^{-6}~m,"

infrared light.